OpenStudy (anonymous):

How many grams of H2O are produced when 12 moles of O2 react? Pleasae explain in detail!

OpenStudy (anonymous):

First you have to get the balanced equation to create water. \[2H_{2} + O_{2}\rightarrow2H _{2}O\] Given that you have 12 moles of O2, all you have to do is do some simple dimensional analysis. \[12 moles O _{2}\times \frac{ 2 moles H _{2} O}{ 1 mole O _{2} }\times \frac{ 18.0194 g H _{2} O }{1 mol }\] Since initially you have 12 moles of O2. All you have to do is to rely on the balanced chemical formula for you to get the mass of H2O. Notice how the units of each compound cancels and only the mass of H2O remains. I'll leave the rest to you. :)